OpenStudy (anonymous):

An object moves according to the position function y =1/(t+2), t>0. Find the velocity as a function of t?? Do I use f(t)-f(a)/t-a or derivative? How?

OpenStudy (anonymous):

take the derivative : y(t)=1/(t+2) -> y=(t+2)^-1 y'(t) = -1*(t+2)^-2 y'(t) = 1/(t+2)^2

OpenStudy (anonymous):

oops -1/(t+2)^2